# Optimal Control of a Subdiffusion model using Dirichlet-Neumann and Neumann-Neumann Waveform Relaxation Algorithms

### v1

@@ Abstract
This paper investigates the convergence behavior of the Dirichlet-Neumann and Neumann-Neumann waveform relaxation algorithms for optimal control problem with sub-diffusion PDE as constraint. The algorithms are applied to regular domains in 1D for multiple subdomains, and the impact of different constant values of the generalized diffusion coefficient on the algorithms’ convergence is analyzed. The convergence rate of the algorithms is analyzed as the fractional order of the time derivative changes.
@@ 1 Optimal Control of a subdiffusion model
Let, $\Omega$ be the open bounded subset in $\mathbb{R}^{d}$ with boundary $\partial\Omega$, and $[0,T]$ is the time domain with $T>0$. First consider the quadratic cost functional
@@
where $u(\boldsymbol{x},t)$ is the control variable and $y_{Q}(\boldsymbol{x},t)$ be the target state. $\sigma$ be the regularization parameter, which mainly use to penalize the control variable. Our objective is to achieve the target state $y_{Q}(\boldsymbol{x},t)$ using the control $u(\boldsymbol{x},t)$, so that the cost functional $J(y,u)$ will be minimum. Here, the state variable $y(\boldsymbol{x},t)$ is governed by the fractional diffusion equation of order $\alpha>0$ given as follows:
@@
(2) $\begin{cases}D^{\alpha}_{RL}y=\nabla\cdot\left(\kappa(\boldsymbol{x},t)\nabla y\right)+f(\boldsymbol{x},t)+u,&\textrm{in}\;\Omega\times(0,T),\\ y(\boldsymbol{x},t)=g(\boldsymbol{x},t),&\textrm{on}\;\partial\Omega\times(0,T),\\ I^{1-\alpha}_{RL}y(\boldsymbol{x},0^{+})=y_{0}(\boldsymbol{x}),&\textrm{in}\;\Omega.\end{cases}$
@@
Here, $\kappa(\boldsymbol{x},t)>0$ is the dimensionless diffusion coefficient and $D^{\alpha}_{RL}$ be the left sided Riemann-Liouville fractional derivative [7] defined for order $\alpha$, $m-1<\alpha<m$ and $m\in\mathbb{N}$ as follows:
@@
And $I^{\alpha}_{RL}$ represents the Riemann-Liouville fractional integral of order $\alpha>0$ on a continuous function $f(t)$ defined as follows:
@@
For $10<\alpha<$ 1 and $f,u\in L^{2}((0,T)\times\Omega)$ the problem (2) has a unique solution $u\in L^{2}((0,T);H^{2}(\Omega)\cap H^{1}_{0}(\Omega))$ for zero Dirichlet boundary condition, see  [5]. And $J$ be the convex function, so unique infimum exist. If $u\in L^{2}((0,T)\times\Omega)$ be the solution of (1) then using the Euler-Lagrange optimality conditions we have adjoint variable $p\in L^{2}((0,T)\times\Omega)$ such that $(u,y,p)$ satisfy the optimality system  [4]:
(3) $\begin{cases}D^{\alpha}_{RL}y=\nabla\cdot\left(\kappa(\boldsymbol{x},t)\nabla y\right)+f(\boldsymbol{x},t)+u,&\textrm{in}\;\Omega\times(0,T),\\ y(\boldsymbol{x},t)=g(\boldsymbol{x},t),&\textrm{on}\;\partial\Omega\times(0,T),\\ I^{1-\alpha}_{+}y(\boldsymbol{x},0)=y_{0}(\boldsymbol{x}),&\textrm{in}\;\Omega.\end{cases}$
(4) $\begin{cases}{}_{T}D^{\alpha}_{t}p=\nabla\cdot\left(\kappa(\boldsymbol{x},t)\nabla p\right)+y-y_{Q},&\textrm{in}\;\Omega\times(0,T),\\ p(\boldsymbol{x},t)=0,&\textrm{on}\;\partial\Omega\times(0,T),\\ p(\boldsymbol{x},T)=0,&\textrm{in}\;\Omega.\end{cases}$
with the condition $\sigma u=-p$. Here ${}_{T}D^{\alpha}_{t}$ be the right sided Caputo fractional derivative [7] defined for order $\alpha$, $m-1<\alpha<m$ and $m\in\mathbb{N}$ as follows:
@@
2 The Dirichlet-Neumann Waveform Relaxation algorithm
The DNWR method [2] is a semi-parallel iterative algorithm, with the exception of the two sub-domain case. This algorithm combines the substructuring DD method for space and the waveform relaxation method for time. Let us consider two sub-domains $\Omega_{1}$ and $\Omega_{2}$ of $\Omega$ that do not overlap, with the interface $\Gamma:=\partial\Omega_{1}\cap\partial\Omega_{2}$. The solution $u$ is restricted to $u_{j}$ on $\Gamma$, where $j=1,2$, and $\boldsymbol{n}_{j}$ represents the unit outward normal on $\Gamma$ for $\Omega{j}$.
An initial guess $\xi^{(0)}(\boldsymbol{x},t)$ for state equation and $\zeta^{(0)}(\boldsymbol{x},t)$ for adjoint equation are selected throughout the interface $\Gamma\times(0,T)$. The solutions $(y_{j}^{(k)},p_{j}^{(k)}),j=1,2,\text{ for }k=1,2,\ldots$ are then obtained through the following Dirichlet-Neumann steps.
(5) $\begin{array}[]{rcll}\begin{cases}D^{\alpha}_{RL}y_{1}^{(k)}=\nabla\cdot\left(\kappa_{1}\nabla y_{1}^{(k)}\right)+f-\frac{p_{1}^{(k)}}{\sigma},&\textrm{in}\;\Omega_{1},\\ I^{1-\alpha}_{RL}y_{1}^{(k)}(\boldsymbol{x},0)=y_{0}(\boldsymbol{x}),&\textrm{in}\;\Omega_{1},\\ y_{1}^{(k)}=\xi^{(k-1)},&\textrm{on}\;\Gamma,\\ y_{1}^{(k)}=g,&\textrm{on}\;\partial\Omega_{1}\setminus\Gamma,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}D^{\alpha}_{RL}y_{2}^{(k)}=\nabla\cdot\left(\kappa_{2}\nabla y_{2}^{(k)}\right)+f-\frac{p_{2}^{(k)}}{\sigma},&\textrm{in}\;\Omega_{2},\\ I^{1-\alpha}_{RL}y_{2}^{(k)}(\boldsymbol{x},0)=y_{0}(\boldsymbol{x}),&\textrm{in}\;\Omega_{2},\\ \kappa_{2}\partial_{\boldsymbol{n}_{2}}y_{2}^{(k)}=-\kappa_{1}\partial_{\boldsymbol{n}_{1}}y_{1}^{(k)},&\textrm{on}\;\Gamma,\\ y_{2}^{(k)}=g,&\textrm{on}\;\partial\Omega_{2}\setminus\Gamma,\end{cases}\end{array}$
and the adjoint problem will be:
(6) $\begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}p_{1}^{(k)}=\nabla\cdot\left(\kappa_{1}\nabla p_{1}^{(k)}\right)+y_{1}^{(k)}-y_{Q}|_{1},&\textrm{in}\;\Omega_{1},\\ p_{1}^{(k)}(\boldsymbol{x},T)=0,&\textrm{in}\;\Omega_{1},\\ p_{1}^{(k)}=\zeta^{(k-1)},&\textrm{on}\;\Gamma,\\ p_{1}^{(k)}=g,&\textrm{on}\;\partial\Omega_{1}\setminus\Gamma,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}p_{2}^{(k)}=\nabla\cdot\left(\kappa_{2}\nabla p_{2}^{(k)}\right)+y_{2}^{(k)}-y_{Q}|_{2},&\textrm{in}\;\Omega_{2},\\ p_{2}^{(k)}(\boldsymbol{x},T)=0,&\textrm{in}\;\Omega_{2},\\ \kappa_{2}\partial_{\boldsymbol{n}_{2}}p_{2}^{(k)}=-\kappa_{1}\partial_{\boldsymbol{n}_{1}}p_{1}^{(k)},&\textrm{on}\;\Gamma,\\ p_{2}^{(k)}=g,&\textrm{on}\;\partial\Omega_{2}\setminus\Gamma,\end{cases}\end{array}$
and then update the interface with the relaxation parameter $\theta\in(0,1]$ by
(7) $\xi^{(k)}(\boldsymbol{x},t)$ $=\theta u_{2}^{(k)}\left|{}_{\Gamma\times(0,T)}\right.+(1-\theta)\xi^{(k-1)}(\boldsymbol{x},t),$
(8) $\zeta^{(k)}(\boldsymbol{x},t)$ $=\phi p_{2}^{(k)}\left|{}_{\Gamma\times(0,T)}\right.+(1-\phi)\zeta^{(k-1)}(\boldsymbol{x},t),$
The primary objective of our analysis is to investigate how the update component error $\omega^{(k-1)}(\boldsymbol{x},t):=y|_{\Gamma\times(0,T)}-\xi^{(k-1)}(\boldsymbol{x},t)$, and $\upsilon^{(k-1)}(\boldsymbol{x},t):=p|_{\Gamma\times(0,T)}-\zeta^{(k-1)}(\boldsymbol{x},t)$, where $u|_{\Gamma\times(0,T)}$ and $p|_{\Gamma\times(0,T)}$ are the exact solution on the interface, converges to zero. Linearity suffices to consider $f(\boldsymbol{x},t)=0$, $g(\boldsymbol{x},t)=0$ for the convergence of error equation with $u_{0}(\boldsymbol{x})=0$ and $\upsilon_{0}(\boldsymbol{x})=0$ in (5) and (6) respectively. This will be studied in section 6.
3 The Neumann-Neumann Waveform Relaxation algorithm
We are considering the NNWR for multiple subdomains setting, where the model problem (3) and (4) on spatial domain $\Omega$ is divided into multiple non-overlapping subdomains $\Omega_{i}$, where $1\leq i\leq N$, and without cross points. We define $\Gamma_{i,j}$ as the intersection of $\partial\Omega_{i}$ and $\partial\Omega_{j}$ for $j=i+1,i-1$. On each subdomain $\Omega_{i}$, we employ the use of unit outward normals $\boldsymbol{n}{i,i-1}$ and $\boldsymbol{n}{i,i+1}$ on the interfaces $\Gamma_{i,i-1}$ and $\Gamma_{i,i+1}$ respectively. Given the initial guess $\xi_{i,j}^{(0)}(\boldsymbol{x},t)$ for state equation and $\zeta_{i,j}^{(0)}(\boldsymbol{x},t)$ for adjoint equation, along the interfaces $\Gamma_{i,j}\times(0,T)$, the NNWR algorithm can be expressed as follows:
(9) $\begin{array}[]{rcll}\begin{cases}D^{\alpha}_{RL}y_{i}^{(k)}=\nabla\cdot\left(\kappa_{i}\nabla y_{i}^{(k)}\right)+f-\frac{p_{i}^{(k)}}{\sigma},&\mbox{in $\Omega_{i}$},\\ I^{1-\alpha}_{RL}y_{i}^{(k)}(\boldsymbol{x},0)=y_{0}(\boldsymbol{x}),&\mbox{in$\Omega_{i}$},\\ y_{i}^{(k)}=g,&\mbox{on $\partial\Omega_{i}\cap\partial\Omega$},\\ y_{i}^{(k)}=\xi_{i,j}^{(k-1)},&\mbox{on $\Gamma_{i,j}$},\end{cases}\end{array}\begin{array}[]{rcll}\begin{cases}D^{\alpha}_{RL}\psi_{i}^{(k)}=\nabla\cdot\left(\kappa_{i}\nabla\psi_{i}^{(k)}\right)-\frac{\chi_{i}^{(k)}}{\sigma},&\mbox{in $\Omega_{i}$},\\ I^{1-\alpha}_{RL}\psi_{i}^{(k)}(\boldsymbol{x},0)=0,&\mbox{in $\Omega_{i}$},\\ \psi_{i}^{(k)}=0,&\mbox{on $\partial\Omega_{i}\cap\partial\Omega$},\\ \kappa_{i}\partial_{\boldsymbol{n}_{i,j}}\psi_{i}^{(k)}=\kappa_{i}\partial_{\boldsymbol{n}_{i,j}}y_{i}^{(k)}+\kappa_{j}\partial_{\boldsymbol{n}_{j,i}}y_{j}^{(k)},&\mbox{on $\Gamma_{i,j}$}.\end{cases}\end{array}$
(10) $\begin{array}[]{rcll}\begin{cases}{}_{T}D^{\alpha}_{t}p_{i}^{(k)}=\nabla\cdot\left(\kappa_{i}\nabla p_{i}^{(k)}\right)+y_{i}^{(k)}-y_{Q}|_{i},&\mbox{in $\Omega_{i}$},\\ p_{i}^{(k)}(\boldsymbol{x},T)=0,&\mbox{in $\Omega_{i}$},\\ p_{i}^{(k)}=0,&\mbox{on $\partial\Omega_{i}\cap\partial\Omega$},\\ p_{i}^{(k)}=\zeta_{i,j}^{(k-1)},&\mbox{on $\Gamma_{i,j}$},\end{cases}\end{array}\begin{array}[]{rcll}\begin{cases}{}_{T}D^{\alpha}_{t}\chi_{i}^{(k)}=\nabla\cdot\left(\kappa_{i}\nabla\chi_{i}^{(k)}\right)+\psi_{i}^{(k)},&\mbox{in $\Omega_{i}$},\\ \chi_{i}^{(k)}(\boldsymbol{x},T)=0,&\mbox{in $\Omega_{i}$},\\ \chi_{i}^{(k)}=0,&\mbox{on $\partial\Omega_{i}\cap\partial\Omega$},\\ \kappa_{i}\partial_{\boldsymbol{n}_{i,j}}\chi_{i}^{(k)}=\kappa_{i}\partial_{\boldsymbol{n}_{i,j}}p_{i}^{(k)}+\kappa_{j}\partial_{\boldsymbol{n}_{j,i}}p_{j}^{(k)},&\mbox{on $\Gamma_{i,j}$}.\end{cases}\end{array}$
The formula is then used to update the interface values:
(11) $\xi_{i,j}^{(k)}(\boldsymbol{x},t)$ $=\xi_{i,j}^{(k-1)}(\boldsymbol{x},t)-\theta_{i,j}\left(\psi_{i}^{(k)}\left|{}_{\Gamma_{i,j}\times(0,T)}\right.+\psi_{j}^{(k)}\left|{}_{\Gamma_{i,j}\times(0,T)}\right.\right),$
(12) $\zeta_{i,j}^{(k)}(\boldsymbol{x},t)$ $=\zeta_{i,j}^{(k-1)}(\boldsymbol{x},t)-\phi_{i,j}\left(\chi_{i}^{(k)}\left|{}_{\Gamma_{i,j}\times(0,T)}\right.+\chi_{j}^{(k)}\left|{}_{\Gamma_{i,j}\times(0,T)}\right.\right),$
Our goal is to study the convergence of error terms on artificial boundary $\omega_{i,j}^{(k-1)}(\boldsymbol{x},t):=y_{i}|_{\Gamma_{i,j}\times(0,T)}-\xi_{i,j}^{(k-1)}(\boldsymbol{x},t)$ and $\upsilon_{i,j}^{(k-1)}(\boldsymbol{x},t):=p_{i}|_{\Gamma_{i,j}\times(0,T)}-\zeta_{i,j}^{(k-1)}(\boldsymbol{x},t)$ as $k\to\infty$.

### v2

# Optimal Control of a Subdiffusion model using Neumann-Neumann and Dirichlet-Neumann Waveform Relaxation Algorithms

@@ Abstract
This article investigates the convergence patterns exhibited by Neumann-Neumann as well as Dirichlet-Neumann waveform relaxation algorithms when applied to distributed optimal control problems subject to sub-diffusive and diffusive partial differential equation (PDE) constraints. Focusing on regular 1D domains with multiple subdomains, we examine the influence of varying values of the diffusion coefficient on algorithms’ convergence. By adopting semi-discrete approach, we discretize in time while maintaining continuity in space. Our simple but elegant approach of a both-sided graded mesh in time simplifies analysis and enhances convergence compared to traditional mesh grading used in time-fractional PDEs.
keywords:
Optimal Control, Domain Decomposition, Sub-diffusion, Neumann-Neumann Waveform Relaxation, Dirichlet-Neumann Waveform Relaxation
1 Introduction
This article is inspired by Gabriele Ciaramella, Laurence Halpern, and Luca Mechelli’s work titled "Convergence Analysis and Optimization of a Robin Schwarz Waveform Relaxation Method for Time-Periodic Parabolic Optimal Control Problems" [ciaramella2024convergence]. In their study, they addressed time-periodic reaction-diffusion partial differential equations as constraints for distributed optimal control problems and employed the Optimized Schwarz waveform relaxation algorithm for solution. In our work, we extend their approach with some modification by employing both Neumann-Neumann and Dirichlet-Neumann waveform relaxation (NNWR, DNWR) algorithms to solve distributed optimal control problems subject to sub-diffusion and diffusion partial differential equation constraints. Furthermore, we relax the time-periodic assumption.
Fractional calculus has become pivotal in modeling complex systems in contemporary research [rossikhin2010application, povstenko2020fractional, qin2017multi, quiroga2015adjoint]. These applications are particularly significant where the current state of evolution depends not only on the immediate previous state but on the entire history, leading to computationally expensive and memory-intensive numerical procedures [qian2017certified]. To tackle these challenges, distributing computations across multiple processors becomes essential. Domain Decomposition (DD) methods [schwarz1870ueber, lions1990schwarz, dryja1990some, cai1999restricted] are highly efficient in this regard. DD methods can be applied in discrete as well as continuous settings. When time evolution is involved, waveform relaxation techniques coupled with DD is introduced [gander1998space, giladi2002space]. DNWR, NNWR, OSWR, and SWR are examples of substructuring waveform relaxation techniques[gander2021dirichlet, etna_vol45_pp424-456, mandal2017neumann, gander2007optimized]. The effectiveness of DNWR and NNWR algorithms for sub-diffusion and diffusion-wave equations has been extensively investigated in [sana2023dirichlet, sana2023dirichlet_a].
The history of optimal control problems spans over three centuries. Modern optimal control theory, as we know it, was developed by L. S. Pontryagin [boltyanskiy1962mathematical]. J. L. Lions played a substantial role in advancing the theory of optimal control of PDEs [lions1971optimal]. For a comprehensive historical survey of optimal control, refer to the works of Sussmann et al. [sussmann1997300] and Fernandez [fernandez2003control]. Solutions to optimal control problems generally fall into two categories: direct and indirect methods. In the direct method, also known as "discretize then optimize," the continuous objective functional and constraints are first converted into discrete algebraic functions and equations, which are then solved using nonlinear programming techniques. In contrast, the indirect method, also known as "optimize then discretize," involves deriving optimality conditions, leading to a coupled system of PDEs, which are subsequently discretized using numerical techniques. If the control variable acts on the boundary, it is termed as boundary control, whereas if it affects the PDE as a forcing condition, the control problem is referred to as distributed control. In this article, we employ the indirect method to solve distributed optimal control problem.
The study of optimal control in continuous cases subject to sub-diffusive PDEs was pioneered by Mophou et.al. [mophou2011optimal_a, mophou2011optimal]. The main themes of their studies include the existence and uniqueness of solutions to control problems and the derivation of coupled PDEs. We apply DNWR and NNWR algorithms to solve these coupled systems. DD methods have proven effective in handling PDE-constrained optimization problems [bensoussan1973methode, benamou1996domain, benamou1997domain].
This article is structured in the following manner: In Section 2, we first consider the distributed optimal control subject to sub-diffusion and diffusion constraints. Then, we use first-order optimality conditions to derive the state and adjoint equations. Sections 3 through 4 present the derivation of the DNWR and NNWR algorithms for coupled PDE systems. In Section 5, we establish some lemmas required for the primary analysis of convergence. Section 6 introduces a both-sided graded mesh. Sections 7 and 8 provide the main convergence results for the DNWR and NNWR, respectively. Finally, in Section 9, we numerically verify all analytical results.
@@ 2 Optimal Control of a subdiffusion model
Let, $\Omega$ be the open bounded subset in $\mathbb{R}^{d}$ with boundary $\partial\Omega$, and time window $[0,T]$ with $T>0$. Let the quadratic cost functional be
@@
where $u(\boldsymbol{x},t)$ is the control variable and $y_{Q}(\boldsymbol{x},t)$ be the target state. $\sigma$ be the regularization parameter, which mainly use to penalize the control variable. Our objective is to achieve the target state $y_{Q}(\boldsymbol{x},t)$ using the control $u(\boldsymbol{x},t)$, so that the cost functional $J(y,u)$ will be minimum. The state variable $y(\boldsymbol{x},t)$ is governed by the fractional diffusion equation of order $\alpha>0$ given as follows:
@@
(2) $\begin{cases}D^{\alpha}_{RL}y(t,\boldsymbol{x})=\nabla\cdot\left(\kappa(t,\boldsymbol{x})\nabla y(t,\boldsymbol{x})\right)+f(t,\boldsymbol{x})+u(t,\boldsymbol{x}),&\textrm{in}\;(0,T)\times\Omega,\\ y(t,\boldsymbol{x})=g(t,\boldsymbol{x}),&\textrm{on}\;(0,T)\times\partial\Omega,\\ I^{1-\alpha}_{RL}y(0^{+},\boldsymbol{x})=y_{0}(\boldsymbol{x}),&\textrm{in}\;\Omega.\end{cases}$
@@
Here, $\kappa(t,\boldsymbol{x})>0$ is the dimensionless diffusion coefficient and $D^{\beta}_{RL}$ be the left sided Riemann-Liouville fractional derivative [podlubny1998fractional] defined for any fractional order $\beta$, $1<\beta<n, nn\in\mathbb{N}$ as follows:
$D^{\beta}_{RL}f(t):=\frac{1}{\Gamma(n-\beta)}\frac{d}{dt^{n}}\int_{0}^{t}(t-\tau)^{n-\beta-1}f(\tau)d\tau.$
And $I^{\beta}_{RL}$ represents the Riemann-Liouville fractional integral of order $\beta>0$ on a continuous function $f(t)$ defined as follows:
$I^{\beta}_{RL}f(t):=\frac{1}{\Gamma(\beta)}\int_{0}^{t}(t-\tau)^{\beta-1}f(\tau)d\tau.$
For $0<\alpha<1$ and $f,u\in L^{2}((0,T)\times\Omega)$ the problem (2) has a unique solution $y\in L^{2}((0,T);H^{2}(\Omega)\cap H^{1}_{0}(\Omega))$ for zero Dirichlet boundary condition, see  [mophou2011optimal]. And the objective functional $J$ is convex, so unique infimum exist. If $y\in L^{2}((0,T)\times\Omega)$ be the solution of (1) then using the Euler-Lagrange optimality conditions we have adjoint variable $p\in L^{2}((0,T)\times\Omega)$ such that $(u,y,p)$ satisfy the optimality system  [mophou2011optimal_a]:
(3) $\begin{cases}D^{\alpha}_{RL}y(t,\boldsymbol{x})=\nabla\cdot\left(\kappa(t,\boldsymbol{x})\nabla y(t,\boldsymbol{x})\right)+f(\boldsymbol{x},t)+u(t,\boldsymbol{x}),&\textrm{in}\;(0,T)\times\Omega,\\ y(t,\boldsymbol{x})=g(t,\boldsymbol{x}),&\textrm{on}\;(0,T)\times\partial\Omega,\\ I^{1-\alpha}_{RL}y(0^{+},\boldsymbol{x})=y_{0}(\boldsymbol{x}),&\textrm{in}\;\Omega.\end{cases}$
(4) $\begin{cases}{}_{T}D^{\alpha}_{t}p(t,\boldsymbol{x})=\nabla\cdot\left(\kappa(t,\boldsymbol{x})\nabla p(t,\boldsymbol{x})\right)+y(t,\boldsymbol{x})-y_{Q}(t,\boldsymbol{x}),&\textrm{in}\;(0,T)\times\Omega,\\ p(t,\boldsymbol{x})=0,&\textrm{on}\;(0,T)\times\partial\Omega,\\ p(T,\boldsymbol{x})=0,&\textrm{in}\;\Omega.\end{cases}$
with the condition $\sigma u=-p$. $_{T}D^{\beta}_{t}$ be the right sided Caputo fractional derivative [podlubny1998fractional] of order $\beta$, $1<\alpha<n, nn\in\mathbb{N}$ defined as follows:
$_{T}D^{\beta}_{t}f(t):=-\frac{1}{\Gamma(n-\beta)}\int_{t}^{T}(\tau-t)^{n-\beta-1}f^{(n)}(\tau)d\tau.$
3 The Dirichlet-Neumann Waveform Relaxation algorithm
The DNWR technique [gander2021dirichlet] is an iterative algorithm that operates semi-parallelly, except in the scenario of two sub-domains. It merges the domain decomposition method for spatial resolution with the waveform relaxation method for temporal resolution. Suppose we have two non-overlapping sub-domains, $\Omega_{1}$ and $\Omega_{2}$, within $\Omega$, with their interface denoted as $\Gamma:=\partial\Omega_{1}\cap\partial\Omega_{2}$. The solution pair $(y,p)$ of state and adjoint equations and diffusion parameter $\kappa(t,\boldsymbol{x})$ in (3)-(4) are restricted to $(y_{j},p_{j})$ and constant $\kappa_{j}$ on $\Omega_{j}$, where $j=1,2$. Let $\boldsymbol{n}_{j}$ denote the unit outward normal vector on $\Gamma$ corresponding to $\Omega{j}$. A starting estimate $\xi^{(0)}(t,\boldsymbol{x})$ for the state equation and $\zeta^{(0)}(t,\boldsymbol{x})$ for the adjoint equation are chosen across the interface $(0,T)\times\Gamma$. The Dirichlet and Neumann steps for state problems yield iterative solutions $y_{j}^{(k)},\text{ where }j=1,2,\text{ and }k=1,2,\ldots$, given by
(5) $\begin{array}[]{rcll}\begin{cases}D^{\alpha}_{RL}y_{1}^{(k)}=\nabla\cdot\left(\kappa_{1}\nabla y_{1}^{(k)}\right)+f-\frac{p_{1}^{(k)}}{\sigma},&\textrm{in}\;\Omega_{1},\\ I^{1-\alpha}_{RL}y_{1}^{(k)}(0,\boldsymbol{x})=y_{0}(\boldsymbol{x}),&\textrm{in}\;\Omega_{1},\\ y_{1}^{(k)}=\xi^{(k-1)},&\textrm{on}\;\Gamma,\\ y_{1}^{(k)}=g,&\textrm{on}\;\partial\Omega_{1}\setminus\Gamma,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}D^{\alpha}_{RL}y_{2}^{(k)}=\nabla\cdot\left(\kappa_{2}\nabla y_{2}^{(k)}\right)+f-\frac{p_{2}^{(k)}}{\sigma},&\textrm{in}\;\Omega_{2},\\ I^{1-\alpha}_{RL}y_{2}^{(k)}(0,\boldsymbol{x})=y_{0}(\boldsymbol{x}),&\textrm{in}\;\Omega_{2},\\ \kappa_{2}\partial_{\boldsymbol{n}_{2}}y_{2}^{(k)}=-\kappa_{1}\partial_{\boldsymbol{n}_{1}}y_{1}^{(k)},&\textrm{on}\;\Gamma,\\ y_{2}^{(k)}=g,&\textrm{on}\;\partial\Omega_{2}\setminus\Gamma,\end{cases}\end{array}$
and for the adjoint problems:
(6) $\begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}p_{1}^{(k)}=\nabla\cdot\left(\kappa_{1}\nabla p_{1}^{(k)}\right)+y_{1}^{(k)}-y_{Q}|_{1},&\textrm{in}\;\Omega_{1},\\ p_{1}^{(k)}(T,\boldsymbol{x})=0,&\textrm{in}\;\Omega_{1},\\ p_{1}^{(k)}=\zeta^{(k-1)},&\textrm{on}\;\Gamma,\\ p_{1}^{(k)}=0,&\textrm{on}\;\partial\Omega_{1}\setminus\Gamma,\end{cases}\end{array}\ \begin{array}[]{rcll}\begin{cases}{}_{T}D_{t}^{\alpha}p_{2}^{(k)}=\nabla\cdot\left(\kappa_{2}\nabla p_{2}^{(k)}\right)+y_{2}^{(k)}-y_{Q}|_{2},&\textrm{in}\;\Omega_{2},\\ p_{2}^{(k)}(T,\boldsymbol{x})=0,&\textrm{in}\;\Omega_{2},\\ \kappa_{2}\partial_{\boldsymbol{n}_{2}}p_{2}^{(k)}=-\kappa_{1}\partial_{\boldsymbol{n}_{1}}p_{1}^{(k)},&\textrm{on}\;\Gamma,\\ p_{2}^{(k)}=0,&\textrm{on}\;\partial\Omega_{2}\setminus\Gamma.\end{cases}\end{array}$
Subsequently, modify the interface after incorporating the relaxation parameters $\theta,\phi\in(0,1)$, by
(7) $\xi^{(k)}(t,\boldsymbol{x})$ $=\theta y_{2}^{(k)}\left|{}_{(0,T)\times\Gamma}\right.+(1-\theta)\xi^{(k-1)}(t,\boldsymbol{x}),$
(8) $\zeta^{(k)}(t,\boldsymbol{x})$ $=\phi p_{2}^{(k)}\left|{}_{(0,T)\times\Gamma}\right.+(1-\phi)\zeta^{(k-1)}(t,\boldsymbol{x}).$
The primary goal of our study is to examine the convergence to zero of the update part errors $\omega^{(k-1)}(t,\boldsymbol{x}):=y|_{(0,T)\times\Gamma}-\xi^{(k-1)}(t,\boldsymbol{x})$ and $\upsilon^{(k-1)}(t,\boldsymbol{x}):=p|_{(0,T)\times\Gamma}-\zeta^{(k-1)}(\boldsymbol{x},t)$, where $y|_{(0,T)\times\Gamma}$ and $p|_{(0,T)\times\Gamma}$ represent the exact interface solution. Linearity allows us to consider $f(t,\boldsymbol{x})=0$, $g(t,\boldsymbol{x})=0$, and $y_{0}(\boldsymbol{x})=0$ for the convergence of the error equations in (5) and (6), respectively. This analysis will be conducted in section 7.
4 The Neumann-Neumann Waveform Relaxation algorithm
We are exploring the NNWR algorithm for a scenario involving multiple subdomains. In this context, the model problem (3)- (4) is defined over the spatial domain $\Omega$, which is partitioned into multiple disjoint subdomains $\Omega_{i}$, where $1\leq i\leq N$, with no cross point. We designate $\Gamma_{i,j}$ as the boundary where $\partial\Omega_{i}$ intersects with $\partial\Omega_{j}$ for j=i1,i+1j=i-1,i+1. Within each subdomain $\Omega_{i}$, we define unit outward normals denoted as $\boldsymbol{n}_{i,i-1}$ and $\boldsymbol{n}_{i,i+1}$ on the boundaries $\Gamma_{i,i-1}$ and $\Gamma_{i,i+1}$, respectively. Given initial guesses $\xi_{i,j}^{(0)}(\boldsymbol{x},t)$ for the state equation and $\zeta_{i,j}^{(0)}(\boldsymbol{x},t)$ for the adjoint equation along the interfaces $(0,T)\times\Gamma_{i,j}$, the NNWR algorithm for optimal control problem can be articulated in the following manner:
@@ 4 The Neumann-Neumann Waveform Relaxation algorithm
Subsequently, the update formula is employed to refresh the interface values to get
@@
[…]
Our objective is now to investigate the convergence of the error equations derived from (9)-(10). This analysis involves considering zero forcing terms and initial conditions, namely $f(t,\boldsymbol{x})=0$ and $y_{0}(\boldsymbol{x})=0$, respectively, as well as imposing zero natural boundary conditions. However, we introduce error components on the artificial interfaces defined as $\omega_{i,j}^{(k-1)}(t,\boldsymbol{x}):=y_{i}|_{(0,T)\times\Gamma_{i,j}}-\xi_{i,j}^{(k-1)}(t,\boldsymbol{x})$ and $\upsilon_{i,j}^{(k-1)}(t,\boldsymbol{x}):=p_{i}|_{(0,T)\times\Gamma_{i,j}}-\zeta_{i,j}^{(k-1)}(t,\boldsymbol{x})$ as $k\to\infty$.